East Texas Seasonings is preparing to build one processing center to serve its four sources of seasonings. The four source locations are at coordinates shown below. Also, the volume from each source is provided. What is the center of gravity?


X-coordinate

Y-coordinate

Volume

Athens, Texas

30

30

150

Beaumont, Texas

20

10

350

Carthage, Texas

10

70

100

Denton, Texas

50

50

400





1
Expert's answer
2021-01-26T03:34:21-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & x & y & Volume \\ \hline
 Athens, Texas & 30 & 30 & 150 \\
 \hdashline
 Beaumont, Texas & 20 &10 & 350 \\
\hdashline
 Carthage, Texas & 10 &70 & 100 \\
\hdashline
 Denton, Texas & 50 &50 & 400 \\
\end{array}"

The coordinates of the center of gravity can be identified by


"x_c =\dfrac{\sum x_iV_i}{\sum V_i}"

"y_c =\dfrac{\sum y_iV_i}{\sum V_i}"

"\sum x_iV_i=30(150)+20(350)+10(100)+50(400)"

"=32500"


"\sum y_iV_i=30(150)+10(350)+70(100)+50(400)"

"=35000"


"\sum V_i=150+350+100+400=1000"


"x_c =\dfrac{32500}{1000}=32.5"

"y_c =\dfrac{35000}{1000}=35"

The weighted center of gravity is located at  "x_c=32.5, y_c=35."



"\sum x_i=30+20+10+50=110"

"\sum y_i=30+10+70+50=160"

The simple center of gravity is located at  "x_s=\dfrac{110}{4}=27.5, y_s=\dfrac{160}{4}=40."